OpenStudy (anonymous):

Find the derivative of the function. y = -1/[(t+6)^6] y '(t) =

OpenStudy (anonymous):

\[-1\div(t+6)^{6}\]

OpenStudy (anonymous):

so, when there is a exponent in the denominator, how do you break the function into outside and inside witht he chain rule?

OpenStudy (anonymous):

rewrite y as y=-1(t+6)^-6

OpenStudy (anonymous):

y'=-1*-6*(t+6)^-7 y'=6/(t+6)^7

OpenStudy (anonymous):

where did the 6 in the numerator come from?

OpenStudy (anonymous):

-6*-1

OpenStudy (anonymous):

negative exponent in expression

OpenStudy (anonymous):

instead of dividing by a positive exponent, i multiplied by a negative exponent

OpenStudy (anonymous):

thanks!